Accelerator Update for October 17, 2008 - October 20, 2008
Friday Oct 17
The day shift began with Operations monitoring the Tevatron (TeV) store 6497 (108.90E30) and the Recycler stash (116.5E10), with the Antiproton Source (Pbar) stacking (40.6E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with the MiniBooNE, NuMI and MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.
Pbar stacking was re-established at 11:24 AM.
MTest turned off at 6:03 PM.
Operations terminated store 6497 at 9:58 PM.
Operations put the TeV into shot setup at 10:13 PM.
The 2 nd Pbar transfer to the Tev was lost at 10:53 PM.
Saturday Oct 18
The midnight shift began with Pbar stacking (76.6E10), and with the MiniBooNE and NuMI experiments taking beam. Shot setup was on hold due to problems with the MI-30 kicker.
The reservoir voltage on the MI-30 Kicker was increased and shot setup resumed at 2:48 AM.
Operations established store 6499 at 4:19 AM with an initial luminosity of 254.80E30.
Operations established beam to MTest at 6:06 AM.
MTest went to open access at 9:00 AM.
At 9:27 AM Booster RF 16 tripped off with an indication of missing 480V 3-phase on the bias supply. The trip would not reset. Operations turned on BRF19 and BRF16 remained off.
Operations terminated store 6499 at 9:26 PM.
Operations put the TeV into shot setup at 9:37 PM.
Operations established store 6500 at 11:19 PM with an initial luminosity of 297.56E30.
Sunday Oct 19
The midnight shift began with Operations monitoring the Tevatron (TeV) store 6500 and the Recycler stash, with the Antiproton Source (Pbar) stacking , with the Linac, Booster, and the Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.
The Tev anode supply tripped off at 8:17 AM, causing the Tev beam to abort. An expert was called in to investigate.
The Crobar chassis for the Tev anode supply was replaced and the supply was back on at 9:33 AM.
Operations put the TeV into shot setup at 10:43 AM.
Operations established store 6501 at 12:20 PM with an initial luminosity of 262.78E30.
MI ramp problems inhibited beam periodically between 1:22 PM and 7:22 PM. In all there was about 2 hours of integrated downtime until MI-50 upper power supply was removed from the loop.
Monday Oct 20
The midnight shift began with Operations monitoring store 6501 and the stash, with Pbar stacking, and with NuMI and BNB taking beam.
Between 12:21 AM and 3:36 AM, the Linac was experiencing problems with all of the low energy gradients going to 0 and the Klystron gradients were oscillating between .7 and 1. The problem was traced to a bad 804 MHz amplifier chassis which had to be replaced.
Operations established beam to MTest at 6:39 AM.
The Plans for Monday
The plans for today are to stack, stash, and store.
Comments and Suggestions
|last modified 10/22/2008 email Fermilab|
|Security, Privacy, Legal|